# FreeTacMan: Robot-free Visuo-Tactile Data Collection System for Contact-rich Manipulation

### v2

@@ V Conclusion
Limitation and future work. While FreeTacMan has demonstrated efficacy across a range of challenging tasks, a few limitations remain. Our system currently depends on an external base station for submillimeter-level localization. To eliminate this dependency, we will develop high-precision visual algorithms for collecting precise visuo-tactile data in the wild. We also plan to extend our data collection system and dataset to multi-fingered dexterous hands, enabling the study of more complex and fine-grained manipulation scenarios.
@@ VI Acknowledgments
We gratefully acknowledge Huijie Wang for helping to develop the demonstration page, and Zherui Qiu for assisting with the supervision of the user study. We also thank Yixuan Pan, Qingwen Bu, Zhuoheng Li, Jisong Cai, Yuxiang Lu, and Ningbin Zhang for their valuable insights and constructive discussions. Our appreciation also goes to Yingxin Wang, Zhirui Zhang, Xinyu Yang, Fengjie Shen, Taoyuan Huang, and Lekai Chen for their assistance during the data collection. Finally, we extend our sincere gratitude to JAKA for their generous support in the collection of our dataset.

### v3

@@ V Conclusion
Limitation and future work. While FreeTacManhas demonstrated efficacy across a range of challenging tasks, a few limitations remain. To eliminate dependency on external base stations for submillimeter-level localization, we will develop high-precision visual algorithms for collecting data in the wild. We also plan to extend our system and dataset to dexterous hands and bimanual long-horizon tasks, facilitating study on finer-grained and more complex scenarios.
@@ VI Acknowledgments
We gratefully acknowledge Huijie Wang for helping to develop the demonstration page, and Zherui Qiu for assisting with the supervision of the user study. We also thank Yixuan Pan, Qingwen Bu, Zhuoheng Li, Jisong Cai, Yuxiang Lu, and Ningbin Zhang for their valuable insights and constructive discussions. Our appreciation also goes to Yingxin Wang, Zhirui Zhang, Xinyu Yang, Fengjie Shen, Taoyuan Huang, and Lekai Chen for their assistance during the data collection. Finally, we extend our sincere gratitude to JAKA for their generous support in the collection of our dataset. This work is in part supported by the JC STEM Lab of Autonomous Intelligent Systems funded by The Hong Kong Jockey Club Charities Trust. This work was also partially supported by the NSFC Grant No.52505029, and the STCSM Grant No.25ZR1401191 and No. 24511103400.